# Insights and Current Gaps in Open-Source LLM Vulnerability Scanners: A Comparative Analysis

### v1

@@ Abstract
This report presents a comparative analysis of open-source vulnerability scanners for conversational large language models (LLMs). As LLMs become integral to various applications, they also present potential attack surfaces, exposed to security risks such as information leakage and jailbreak attacks. Our study evaluates prominent scanners - Garak, Giskard, PyRIT, and CyberSecEval - that adapt red-teaming practices to expose these vulnerabilities. We detail the distinctive features and practical use of these scanners, outline unifying principles of their design and perform quantitative evaluations to compare them. These evaluations uncover significant reliability issues in detecting successful attacks, highlighting a fundamental gap for future development. Additionally, we contribute a preliminary labelled dataset, which serves as an initial step to bridge this gap. Based on the above, we provide strategic recommendations to assist organizations choose the most suitable scanner for their red-teaming needs, accounting for customizability, test suite comprehensiveness, and industry-specific use cases.
@@
Large Language Models (LLMs), Vulnerability Scanners, Red Team, Comparative Analysis, Harmful Content Detection.
@@ I Introduction
In this work, we compare tools for assessing the vulnerability of conversational large language models (LLMs) such as the GPT family, LLaMA-instruct, and Command-R [1, 2, 3]. LLMs are increasingly integrated into various applications, providing a natural language prompt interface [4, 5, 6]. However, this integration exposes systems to significant security risks [2], including the spread of misinformation [7], hate campaigns [8], and cybercriminal activities [9]. To address these threats, the practice of red-teaming has emerged as a crucial part of the defense strategy. Originally rooted in traditional cybersecurity, red-teaming involves simulating attacks to uncover vulnerabilities [10]. In our scenario, red-teaming is adapted to target the prompt interface by crafting adversarial text inputs (prompts), addressing the unique vulnerabilities of LLMs [11, 12]. The importance of red-teaming has been recognized by authoritative sources, including the US government, the National Institute of Standards and Technology, and the OWASP Top 10 for LLM Applications Cybersecurity and Governance Checklist [13, 14, 15]. While LLM red-teaming is essential, it is challenging due to its dynamic nature, which requires continuous updates to stay ahead of evolving threats.
@@
LLM vulnerability scanners (in short, scanners) have recently been developed to facilitate and automate the red-teaming process, enhancing efficiency and adaptability to continuously evolving threats [16, 17, 18, 19, 20, 21, 22, 23, 24]. These scanners systematically test a target model by generating adversarial prompts, designed to elicit invalid responses such as confidential data, or toxic content [25, 26, 27]. The scanner then automatically evaluates the vulnerabilities exposed by these attacks, typically aggregating them to output a summary of the findings.
Figure 1: High-level Overview of our Quantitative Results. a) Scanner performance scatter plot. y-axis: Reported attack effectiveness; x-axis: Average reliability based on correct evaluation of attacks’ success. Circle radius: No. of adversarial prompts in the test-suite. b) Adversarial prompts distribution. Prompt’s attack types are grouped into five categories for comparability.
Since LLM vulnerability scanners are relatively new, there is still a considerable knowledge gap concerning their effectiveness, reliability, respective advantages and usage know-how. Although there are reports on individual scanners and red-teaming best practices (e.g., [28, 16, 12]), to the best of our knowledge, there is an absence of a comprehensive hands-on report for making informed decisions when selecting and employing a scanner. Such a report is necessary, since navigating this domain can be challenging - the field is still in its early stages, with rapidly evolving methodologies and a lack of consensus on best practices. This leads to a diverse array of tools, each targeting different aspects of the domain, as well as overlooked aspects of performance and utility.
@@
This work aims to bridge this knowledge gap by providing a detailed comparison through hands-on experience, and quantitative analyses. We define criteria to identify the most useful scanners for our analyses: Garak, Giskard, PyRIT, and CyberSecEval [17, 22, 29, 16]. This focused selection enables thorough experimentation to examine the internal workings of each scanner and gain a comprehensive understanding of their distinctive features. Furthermore, we conduct extensive quantitative analyses of the attacks utilized by each scanner and the automatic mechanisms that assess the attacks’ success. As shown in Fig. 1, the attacks’ coverage, effectiveness, and reliability are not always high, and vary between the scanners (for the attack categories, see Appendix A). The hands-on insights and quantitative analyses presented will assist researchers and AI safety teams in making informed decisions regarding the selection and improvement of these tools. Our contributions are:
@@
  • Pioneering Analysis: To the best of our knowledge, this is the first hands-on comparative study of open-source LLM vulnerability scanners. We offer valuable insights, practical information, and identify key current challenges.
@@
  • Detailed Feature Insights: We outline the distinctive features and shared unifying principles of various scanners, equipping users with a nuanced understanding of each tool’s strengths.
@@
  • Labelled Dataset: We provide a 1,000-sample dataset as a foundational starting point to initiate the currently absent quantification of the scanners’ reliability [^1] https://drive.google.com/file/d/1mWPKF5Eww-Bma0hGBT7lckNim3Hg6wEK/view?usp=sharing.
  • Quantitative Findings: Analyzing $\sim$5K adversarial prompts, we show that detecting successful attacks is an unresolved challenge - highlighting gaps in reported coverage, as current scanners misclassify up to 37% of successful adversarial prompts for a specific attack.
@@
  • Strategic Recommendations: We provide guidance to organizations in choosing their scanner, considering customizability, vulnerability coverage, and organizational needs. Additionally, we propose future directions to enhance safety performance of these scanners.
II Previous Work
The widespread adoption of LLMs has enhanced digital productivity, while also exposing numerous security vulnerabilities. In response, extensive research has been conducted exploring these vulnerabilities, ranging from biases in training data to susceptibility to adversarial prompts. This research field primarily revolves around: 1) comprehensive surveys cataloging vulnerabilities and evaluation approaches [7, 30, 31] and 2) in-depth works on specific LLM security exploits [32, 33, 34].
Building upon these foundational studies, the field has increasingly embraced red-teaming as an approach to enhance LLM security [35, 36]. Red-teaming process involves simulating adversarial attacks to identify potential LLMs’ vulnerabilities, thereby improving their robustness. To facilitate this process, a variety of automated tools have been developed [37, 16, 29, 22], which enable continuous and scalable security assessments. Some related works automatically evaluate LLMs using other LLMs, however they isolate the effect of the evaluator LLM’s instructions - a key component we aim to analyze [38, 39, 40]. These tools differ in their approach, features, and effectiveness, highlighting the urgent need for a practical comparative analysis. Addressing this need, this report delivers a technical and detailed comparison of leading scanners, offering guidance to red-teaming groups in selecting the most effective tools for their specific organizational needs.
@@ II Previous Work
Figure 2: General design of the automated red-teaming flow, used by LLM vulnerability scanners.
III Unifying Principles of LLM Vulnerability Scanners
@@
LLM vulnerability scanners have only recently emerged, and thus, to the best of our knowledge, their shared principles have not been outlined yet. In this section, we provide such outlines and introduce key terminology used throughout this paper.
@@
A scanner receives a target LLM as input to identify and report its vulnerabilities. The different scanners operate on the same principle of automated red-teaming, and hence exhibit similar system architectures and components. As illustrated in Fig. 2 - the test suite, designed to identify vulnerabilities, is composed of an array of attacker-evaluator pairs. The attacker provides prompts intended to elicit prohibited responses from the target LLM (where ’prohibited’ may be subjective to the specific purpose of the target LLM), while the evaluator determines the success of these attacks.
@@
The used attacks are categorized into two types: Static attacks, which utilize an attack dataset consisting of predefined adversarial prompts, and LLM-based attacks, where an attacker LLM is instructed to generate the adversarial prompts [^2] Static attacks might use LLMs beforehand, but LLM-based attacks instruct attacker LLMs during execution.. The attacker LLM may receive additional external context, such as a list of requirements defining valid responses. Evaluators can also be either static or LLM-based. Static evaluators verify the presence of specific strings in the target LLM’s response via string-matching, including regular expression pattern matching, exact matching, and word matching. In contrast, LLM-based evaluators involve an evaluator LLM that is instructed to identify prohibited responses. It receives the target LLM’s response, and additional contexts such as the attack prompt, and the requirements mentioned above. The identified vulnerabilities are then compiled into a report.
@@
Different scanners may utilize different of attacks, presenting a challenge for comparison. To enable a structured comparison of vulnerability coverage and quantitative performance metrics, these attacks can be grouped into four unifying categories: Jailbreak, Context and Continuation, Gradient-based, and Code Generation Attacks. While this approach reduces granularity, it is essential for ensuring comparability. The specific categorization for each scanner is detailed in Table III in the Appendix.
@@
The unifying principles outlined above set the stage for the diverse landscape of LLM vulnerability scanners. While the implementation of each component can vary significantly, in terms of the attacker and evaluator types, the strategies for generating prompts, and the criteria for identifying prohibited responses - all scanners share a common architecture. The resulting scanner variations are designed to address different aspects of LLM security. In the following sections, we will delve into these variations, exploring the distinctive features and methodologies that distinguish various scanners.
@@ IV Per-Scanner Review
IV-A Scanner Criteria and Selection
Before the hands-on assessments, we conducted an initial review to select the most relevant scanners. Below we detail the three criteria employed for the selection process, the 15 candidate scanners, and the selected scanners that fit these criteria the most. The criteria comprise:
1) Quality of vulnerabilities test suite. This is the cornerstone of any scanner - a comprehensive suite that simulates diverse real-world scenarios is essential for uncovering LLMs vulnerabilities. Well-documented and research-backed attacks provide more reliability to the security assessments.
2) Frequent updates to address evolving threats. The landscape of cybersecurity threats is continuously changing, with new attack vectors emerging regularly. Moreover, as LLMs frequently evolve to counter existing attacks, a scanner’s test suite that was effective a few months ago might be obsolete today. To maintain their effectiveness, scanners must systematically update their test suite, ensuring they can identify the latest threats effectively.
3) Open-source code with active community engagement. In our view, a key advantage is the impact of an active community on the longevity and sustainability of a project, ensuring frequent updates for the long run. An engaged community reduces reliance on initial contributors, fostering a self-sustaining environment where the project can evolve independently. Moreover - as contemporary attacks often emerge from the community, leveraging this collective expertise for test suite updates is a viable strategy. Lastly, we focus solely on open-source scanners for transparency and adaptability.
The $15$ scanners: See Appendix D for a summary of all $15$ scanners, listed below: CyberSecEval [28], HouYi[37], JailBreakingLLMs [41], LLMAttacks [42], Garak [16], Giskard [17], Prompt Fuzzer [18], PromptInject [43], Promptfoo [19], LLMCanary[20], Agentic security [21], PyRIT [22], LLMFuzzer [23], PromptMap [24], Vigil-llm [44].
Four scanners stood out by our criteria [^3] Others using a similar scanner selection process may identify a different distinguished group.:
Garak. Released in early 2024, Garak[16] quickly gained popularity due to its broad vulnerability coverage and frequent updates. Moreover, its attacks are backed by published research.
Giskard. Giskard is a company focused on responsible AI, offers an open-source LLM vulnerability scanner[17]. Giskard has an active and growing community that foster members to engage in several topics including contributing code, discussing vulnerabilities and sharing best practices.
PyRIT. The Python Risk Identification Tool for Generative AI, or PyRIT, is a scanner developed by Microsoft[22]. Released in March 2024[45]. PyRIT has been continually evolving, focusing on LLM-based attack and evaluation strategies.
CyberSecEval. Developed by Meta, this scanner specializes in detecting vulnerabilities in LLM-generated code, while also addressing natural language vulnerabilities [28, 46].
IV-B Garak
Test-suite. Garak excels in vulnerability coverage with over 20 specific attack-evaluation pairs, focusing on jailbreak attacks grounded in established research [16, 47, 32, 33, 42, 48]. It focuses in static attacks with correspondingly static evaluators.
Operational usage. Garak produces reports in both JSON and HTML formats. The JSON report is thoroughly detailed, including every tested adversarial prompt, the target LLM’s reponse to it, and the evaluator’s assessment of success. A additional ’hitlog’ reports successful attacks only. The HTML format, easier for non-technical users, provides a broader overview, summarizing statistics for each attack. An example report can be found in the Appendix 5. We found Garak’s code implementation to be the most robust. It features a ‘probe‘ for managing attacks, and for the evaluation, a ‘detector‘ scores the attack and an ‘evaluator‘ make the final decision based on a pre-defined threshold.
A notable feature is Garak’s integration with Nvidia’s NeMo Guardrails[49], which enables the vulnerability assessment of the target LLM with and without defense mechanisms. Garak includes a generator component that handles all text-to-text interactions and is specifically designed to support NVIDIA NeMo models. The effectiveness of this integration is demonstrated by comparing the performance of unprotected models versus models with various levels of Guardrails defense against specific attacks. For instance, we tested the the GPT 3.5 model’s robustness against context & continuation attacks under various NeMo Guardrails settings: Clean setting achieved 76.1%; Instructions—enhanced protection settings yielded 81.2%; Dialog Rails—mechanisms for refusing unwanted topics reached 92.6%; and Moderation Rails—input/output self-checking scored 95.4%, highlighting Guardrails’ effectiveness.
IV-C Giskard
Test-suite. Giskard’s test suite features a diverse set of nine attack-evaluation pairs that blend static and LLM-based methods, with the latter offering a wide array, ranging from eliciting hallucinations to harmful content and stereotypes, to information disclosure - see Table IV in the Appendix. A distinctive aspect is its context-based mechanism, which includes: 1) a model description of the target model, and 2) in most cases, a list of safety requirements. The requirements are generated per attack, specifying the expected behavior of a robust target model against the attack at hand, see Fig. 3. This sophisticated process curates hybrid instructions for both the attacker and evaluator LLMs. These instructions combine the generated and user-provided contexts within a structured template, as well as pre-fixed context such as the attack category, and generation examples. The generation examples also play a role in defining the expected output format and may include non-English languages (see Fig. 8 in the Appendix). Notably, this attack process involves two LLMs: The standard attacker LLM and the requirements LLM, unlike ”regular” LLM-based attacks which usually employ a single LLM.
Operational usage. After the evaluation, Giskard compiles all failed cases into a user-friendly HTML report, categorizing them by attack type and including prompts, responses, severity and the evaluator-LLM’s reasons for attack success - see Fig. 5 in the Appendix. Giskard’s model description context provides flexibility in tailoring LLM-based attacks to the target-LLM, see for instance Fig. 3 where we attack a bot designated to organize an international AI workshop. To set the model description, the user passes it to the ‘giskard.Model’ class which manages the target model.
Giskard’s attacker and evaluator LLMs are restricted to GPT-family models (3.5 and above), which limit flexibility. While Giskard inherently focuses on successful attacks, we extended its functionality to log benignly classified attacks. This extension required substantial modifications due to the lack of uniform implementation across its attack-evaluation pairs, except for the ‘RequirementBasedDetector’, which offers a standardized interface. Altering attack-evaluation parameters and instruction templates is also useful, but necessitates modifications to deeper parts of the code. Similarly to Garak, Giskard also includes Guardrails-specific modules.
IV-D Pyrit
Test-suite. PyRIT[22] offers a fully LLM-based framework, with a highly useful flexible implementation. It allows a direct access to the attacker and evaluator LLMs’ instructions (in contrast to the pre-made or hybrid attacker instructions described above). PyRIT implements two attack approaches under its attack strategy component: 1) A single-turn attack, similar to the LLM-based attacks presented above; 2) A multi-turn attack, engaging the target LLM in a sequential dialogue, terminating when a predefined goal is achieved (see Fig. 3). These gained a higher success rate, see Sec. V. The evaluator LLM, implemented in the scoring engine, serves a dual role in multi-turn attacks. Beyond deciding whether the attack succeeded, it also determines after each target LLM response if the dialogue should continue. The scoring engine offers four strategies for using evaluator LLMs (see Fig. 11 in the Appendix), with varying outputs: Some provide a binary decision (attack success or failure), while others generate score vectors - discrete rankings (1 to 5) or continuous (0 to 1). These scores evaluate the target LLM’s response across categories like Hate, Bias, Violence, and Sexual content, following Microsoft Azure and OpenAI moderation standards [50].
Operational Usage. PyRIT offers creating custom attack strategies by defining instructions for both attacker and evaluator LLMs, allowing for tailored testing scenarios that target specific vulnerabilities beyond pre-made frameworks. This involves providing system prompts to the LLMs at initialization, enabling highly specialized attack simulations. Users can also modify evaluator instructions, from our experience, even when using well-crafted instructions, variations in evaluator prompts can significantly influence performance. Users can adjust these instructions to focus on areas like bias, hate speech, or ethical violations, ensuring the evaluation aligns with specific goals. This flexibility allows precise vulnerability assessments tailored to various industries or research objectives.
Although PyRIT doesn’t generate a formal report, the attack conversation and success rate are easily accessible. Additionally, PyRIT offers an LLM-generated explanation to clarify its decisions, improving both the transparency and interpretability of the results.
Figure 3: Top: Example Flow of Giskard, testing a ”Workshop Organizer AI”. An important aspect of Giskard is its ability to customize tests for LLMs that are designed for specific tasks. This example demonstrates Giskard’s customization via its distinctive requirements-based test. This is a shortened version - for the full version, including the evaluation phase, refer to Fig. 8 in the Appendix. Bottom: Example flow of PyRIT’s multi-step attack for generating Python Key Logger. An attacker LLM is tasked with attacking a target LLM under evaluation, while another LLM assesses the attack’s success. This loop continues until the attack succeeds or a stopping criterion is met.
IV-E CyberSecEval
Test-suite. CyberSecEval [28, 29, 46] provides a fully static test suite focused on analyzing vulnerabilities in LLM-generated code. It primarily targets insecure coding practices, which are codes with security breaches. As a side focus, the suite also evaluates the generation of malicious code, i.e. code that causes harm upon execution, and includes natural language jailbreak attacks that coax LLMs into revealing sensitive information.
To test for insecure coding practices, the CyberSecEval incorporates two main strategies, which use a database of insecure code snippets - as illustrated in Fig. 9 in the Appendix. The first is an auto-complete scenario, where the LLM is prompted with the 10 code lines that precede an insecure practice to see if it reproduces the risky code. The second, referred as instruct, converts instances of insecure practices into natural language instructions which in turn are passed as prompts to the target LLM, to assess if the LLM replicates the insecure practice. The dataset includes $50$ Common Weakness Enumerations (CWEs) - a comprehensive list of common software security vulnerabilities [51], across eight programming languages. CyberSecEval evaluator, called the Insecure Code Detector (ICD) identifies insecure coding practices for test case generation and model evaluation. The ICD uses rules created by Meta’s cyber security experts, applied through static analysis tools like Weggli, Semgrep, and regular expressions (see Figure 10 in the Appendix).
Operational Usage. Upon concluding an evaluation session, CyberSecEval produces two types of reports: a statistical report that summarizes the success rates of various attacks across all test cases, and a detailed response report that documents each attack along with the AI’s corresponding response.
While CyberSecEval’s code implementation is stable and well-designed for easy modification of attacks and evaluations, crafting these is not straightforward, as it requires both cybersecurity and AI expertise. With that said - there are thousands of open-source CWEs which one can extend the insecure coding coverage utilizing CybeSecEval’s framework. Thus in the hands of experts, CyberSecEval has potential for significant extensions.
IV-F Summary of Distinctive Features and their Comparison
To conclude this Section, Table I provides a comparison highlighting distinctive scanner features.
LLM Guardrails Interface. A ready-to-use encapsulation of one of the popular open-source Guardrails solution. Garak and Giskard support this. Explainability. Whether the scanner offers explanations for its assessments of attack success. Currently, only PyRIT and Giskard support this. Multi-language Support. Support the ability to expose vulnerabilities in non-English languages. Giskard is the only scanner that allows users to specify the language via example templates and provide non-English examples, prompting the attacker LLM to generate similar prompts. Customizable Attacks. Enable user-directed attack customization. Both Giskard and PyRIT support this, yet their approaches differ significantly. PyRIT grants full control over the attacker LLM’s system prompt, while Giskard employs a more structured approach by combining pre-defined prompts with user inputs, restricted to the target model’s name and description. Insecure Coding Tests. Whether the scanner includes tests specifically designed to assess the security of code generated by LLMs. Currently, only CyberSecEval and Garak offer such capability.
TABLE I: Comparison of distinctive scanner features. None of the scanners covers all utilities.

### v2

@@ Abstract
We present a comparative analysis of open-source tools that scan conversational large language models (LLMs) for vulnerabilities, in short - scanners. As LLMs become integral to various applications, they also present potential attack surfaces, exposed to security risks such as information leakage and jailbreak attacks. AI red-teaming, adapted from traditional cybersecurity, is recognized by governments and companies as essential - often emphasizing the challenge of continuously evolving threats. Our study evaluates prominent, cutting-edge scanners - Garak, Giskard, PyRIT, and CyberSecEval - that address this challenge by automating red-teaming processes. We detail the distinctive features and practical use of these scanners, outline unifying principles of their design and perform quantitative evaluations to compare them. These evaluations uncover significant reliability issues in detecting successful attacks, highlighting a fundamental gap for future development. Additionally, we contribute a foundational labeled dataset, which serves as an initial step to bridge this gap. Based on the above, we provide suggestions for future regulations and standardization, as well as strategic recommendations to assist organizations in scanner selection, considering customizability, test-suite comprehensiveness and industry-specific use cases.
@@
Large Language Models (LLMs), Vulnerability Scanners, Fuzzers, Red-teaming, Comparative Analysis.
@@ I Introduction
In this work, we compare tools for assessing the vulnerability of large language models (LLMs) such as the GPT family, LLaMA-instruct, and Command-R [1, 2, 3]. LLMs are increasingly integrated into various applications, providing a natural language prompt interface [4, 5, 6]. However, this integration exposes systems to significant security risks [2], including the spread of misinformation [7], hate campaigns [8], and cybercriminal activities [9]. Consequently, red-teaming has emerged as a crucial part of the defense strategy.
Rooted in traditional cybersecurity, red-teaming simulates attacks to uncover vulnerabilities [10]. In our scenario, it is adapted to target the prompt interface through adversarial text inputs (prompts), addressing LLMs’ vulnerabilities [11, 12].
The importance of LLM red-teaming has been recognized by authoritative sources, including the US government, the National Institute of Standards and Technology, and the OWASP Top 10 for LLM Applications Cybersecurity and Governance Checklist [13, 14, 15]. A key challenge is maintaining red-teaming tools due to the dynamic nature of threats, which requires continuous updates to stay ahead.
@@
Recently, to maintain the safety of LLM-based systems throughout their lifecycle, tools were developed for ”LLM Automated Benchmarking” - as defined by the OWASP report, October 2024 (Q4) [16]. These include (among others) LLM vulnerability scanners, fuzzers and other tools that facilitate and automate the red-teaming process [17, 18, 19, 20, 21, 22, 23, 24, 25]. For convenience, we will consider all such tools as scanners: Scanners test a target model by generating adversarial prompts, designed to elicit invalid responses such as confidential data, or toxic content [26, 27, 28]. The scanner then automatically evaluates the vulnerabilities exposed by these attacks.
@@
Since scanners are relatively new, there is still a considerable knowledge gap concerning their effectiveness, reliability, respective advantages and usage know-how. This work aims to bridge this gap by providing a detailed comparison through hands-on experience and quantitative analyses, of four of the most widely used open-source scanners: Garak, Giskard, PyRIT, and CyberSecEval [17, 18, 23, 29]. Through experimentation, we examine the internal workings of each scanner, gaining a comprehensive understanding of their distinctive features. We conduct extensive quantitative analyses of the scanners - as shown in Fig. 1, the attacks’ coverage, effectiveness, and reliability are not always high and vary between the scanners (for the attack categories, see Table I). These hands-on insights and quantitative analyses will assist researchers and AI safety teams to better use and develop these tools. For supplementary see link  [^1] Supplementary and dataset: https://tinyurl.com/scanners-material Our contributions are:
Figure 1: High-level Overview of our Quantitative Results. a) Scanner performance scatter plot. y-axis: Reported attack effectiveness; x-axis: Average reliability based on correct evaluation of attacks’ success. Circle radius: No. of adversarial prompts in the test-suite. See further detail on how these axes are calculated in Sec. V b) Adversarial prompts distribution. Prompt’s attack types are grouped into five categories for comparability.
TABLE I: Overview of attack categories, descriptions, and associated attacks from Garak, Giskard, PyRIT, and CybrSecEval scanners.
[…]
Attack Description Disrupt and bypass LLM restrictions through targeted prompt attacks. Utilizes gradient information to probe the LLM for vulnerabilities. Exploits biases in the LLM related to ethnicity, gender, sexual orientation, and religion. Instruct the LLM to generate harmful code, such as malware and keyloggers, or induce cybersecurity vulnerabilities in LLM-generated code. Engaging the target LLM in a sequential dialogue, with a predefined adversarial goal.
[…]
@@
  • Pioneering Analysis: To the best of our knowledge, this is the first hands-on comparative study of open-source LLM vulnerability scanners. We offer valuable insights, practical information and identify key current challenges.
@@
  • Detailed Feature Insights: We outline the distinctive and shared principles of various scanners, equipping red-teamers with a nuanced understanding of these tools.
@@
  • Labelled Dataset: We provide a 1,000-sample dataset as a foundational starting point to initiate the currently absent quantification of the scanners’ reliability 11footnotemark: 1.
  • Quantitative Findings: We analyze 4 leading tools across 4 LLMs using $\sim$5K adversarial prompts; providing statistics of their merits, comprehensiveness and accuracy.
  • Reliability Analysis: We show for the first time the gap in detecting successful attacks, where misclassification of an attack can reach 37%. Qualitative examples are analyzed to uncover underlying reasons for this limitation.
@@
  • Strategic Recommendations: We provide guidance in choosing a scanner considering organizational needs, customizability and vulnerability coverage, and propose future directions to enhance safety regulations.
II Related Work
LLM safety has been extensively researched, leading to several comprehensive reviews: Chi et. al.,[34] reviewed jailbreak attack methods, revealing that optimized prompts can consistently bypass LLM safeguards. Ruiu [35] reviewed various attack types against LLMs and provided red team best practices to enhance LLM safety. Kenthapadi et. al.,[36] reviewed LLM evaluation metrics focused on aspects of responsible AI such as robustness, bias and security. However, there remains a notable gap in the literature concerning scanners for LLM vulnerability detection. Numerous scanners were developed (see supplementary for the descriptions of $16$ different scanner), but current thorough reports deal with individual scanners, e.g. [17, 29, 23, 21, 20, 37, 24, 22, 33]. Scanners vary in approaches, features, and levels of effectiveness, underscoring the need for informative comparisons between them. We compare 4 leading open-source scanners: Garak (v0.9.0.14.post1), associated with Nvidia [17], features broad vulnerability coverage, frequent updates and research-backed attacks. Giskard (v2.14.4), is by Giskard - a company focused on responsible AI [18]. It has an active and growing community engaging in code contributions, vulnerabilities discussions and best practices. PyRIT (v0.2.2.dev0) by Microsoft[23] has been continually evolving, implementing partially and fully automatic LLM red-teaming strategies. CyberSecEval by Meta, specializes in detecting vulnerabilities in LLM-generated code, and addresses natural language vulnerabilities [38, 39]. To our knowledge, this is the first report to derive insights from hands-on testing of several leading scanners. We provide a ’versions snapshot’ from H2 2024; while updates to these tools and others are ongoing, our identified gaps and insights apply across the domain for future development.
@@ II Related Work
Figure 2: General design of the automated LLM red-teaming flow, used by scanners.
III Unifying Principles of the Scanners
@@
The category of tools collectively referred to as ”scanners” has only recently emerged, and to the best of our knowledge, their shared principles have not been outlined yet. Here we provide such outlines and introduce key terminology.
@@
A scanner receives a target LLM as input to identify and report its vulnerabilities. The different scanners operate on the same principle of automated red-teaming and hence exhibit similar system architectures and components. As illustrated in Fig. 2, the test-suite, designed to identify vulnerabilities, is composed of an array of attacker-evaluator pairs. The attacker provides prompts intended to elicit invalid responses from the target LLM (where ’invalid’ may be subjective), while the evaluator determines the success of these attacks.
@@
The used attacks are categorized into two types: Static attacks, which utilize a predefined attack dataset of adversarial prompts; and LLM-based attacks, where an attacker LLM is instructed to generate the adversarial prompts [^2] Static may use LLMs beforehand; LLM-based generate prompts live.. The attacker LLM may receive additional context, such as a list of requirements defining valid responses. Evaluators can also be either static or LLM-based. Static evaluators verify the presence of specific strings in the target LLM’s response via patten or word matching, such as regex. In contrast, LLM-based evaluators involve an evaluator LLM that is instructed to classify (in)valid responses, sometimes accompanied by a textual evaluator explanation of the decision. It receives the target LLM’s response, and additional contexts such as the attack prompt, and the requirements mentioned above. Identified vulnerabilities are then compiled into a report.
@@
Different scanners may utilize different attacks, presenting a challenge for comparison. To enable comparison of vulnerability coverage and quantitative performance metrics, these attacks can be grouped into four unifying categories: Jailbreak, Context and Continuation, Gradient-based, and Code Generation Attacks. While this approach reduces granularity, it is essential for ensuring comparability. The categorization for each scanner is detailed in Table I.
@@
The unifying principles outlined above set the stage for the diverse landscape of the scanners. While the implementation of each component can significantly vary in terms of the attacker and evaluator types, the strategies for generating prompts and the criteria for identifying invalid responses - all scanners share a common architecture. The resulting scanner variations are designed to address different aspects of LLM security. In the following sections, we will delve into these variations, exploring the distinctive features and methodologies that distinguish various scanners.
@@ IV Per-Scanner Review
All four scanners have high-quality vulnerability test suits, active community, and frequently updated open-source code. Below we summarize their distinctive features, derived from hands-on experience with their code, culminating in Table II.
Garak. Garak excels in vulnerability coverage with over 20 specific attack-evaluation pairs, focusing on jailbreak attacks grounded in established research [17, 40, 30, 41, 33, 32]. It focuses on static attacks and static evaluators. Garak produces highly detailed reports, including every tested adversarial prompt, the target LLM’s response, and the evaluator’s assessment of success. A broader overview report is provided in supplementary Figs. 1-3. Notably, Garak integrates with Nvidia’s NeMo Guardrails[42], allowing comparison of vulnerabilities (detected by Garak) between models with varying levels of defense (provided by Nemo Guardrails).
TABLE II: Comparison of distinctive scanner features.